OpenStudy (anonymous):

Plz Help! What is the exact area of a circle whose diameter is 24 cm?

OpenStudy (unknownunknown):

So Rayne, if the area of a circle is given by: \[\pi r^{2}\] and we know the diameter is 2r. Can you work out its area now?

OpenStudy (anonymous):

OK give me one min.

OpenStudy (unknownunknown):

Sure, take your time.

OpenStudy (anonymous):

I can get it

OpenStudy (unknownunknown):

So what is the answer?

OpenStudy (anonymous):

i ment cant

OpenStudy (anonymous):

I think i am over thinking it

OpenStudy (unknownunknown):

So remember the formula is \[\pi r^{2}\], and we know that d = 24cm. We also know that d = 2r. Solving for r, r = d/2. 24/2 = 12cm. Therefore, r = 12cm, correct? Now, plug that into the equation \[\pi r^{2}\] \[\pi * 12^{2} = ?\]

OpenStudy (anonymous):

Oh I know what I am doing wrong I was doing 24 instead of 12 oops now I maybe I can do it. lol

OpenStudy (anonymous):

114

OpenStudy (anonymous):

right?

OpenStudy (unknownunknown):

Unfortunately not. So, what is 12 * 12?